OpenStudy (anonymous):

2 Resistors of resistance, R = (100 +- 3) Ohm, R2 = (200+-4)Ohm are connected in Parallel and Series. Find the absolute error and % error in the measurement of equivalent resistance are in each combinations I.e Series and Parallel.. ?? PLEASE HELP!!

OpenStudy (anonymous):

Help!!

ganeshie8 (ganeshie8):

lets work series first

ganeshie8 (ganeshie8):

do u knw formula for equivalent resistance for resistors in series ?

OpenStudy (anonymous):

R= R1+r2+r3... and for parallel.. R= R.R1/R1+r2

ganeshie8 (ganeshie8):

good we can start from there : when the 2 resistors are connected in series ideally we expect to see 100 + 200 = 300 ohm. but since these are not ideal, first resistor can change +- 3, that is, it can be anywhere between [97, 103] second resistor can change +- 4, that is, it can be anywhere between [196, 204] so overall, the series connection can see an absolute error of +- 7 ohm and a % error of 7/300 * 100 = 2.33 %

ganeshie8 (ganeshie8):

see if the above make sense, we can do the parallel same way

OpenStudy (anonymous):

thank you.. ^_^

ganeshie8 (ganeshie8):

yw ! try the parallel circuit also... ping me if u get stuck.. .

OpenStudy (anonymous):

okay.. :D thanks alot.. ^_^

OpenStudy (anonymous):

can you just do it for parallel too? ^_^ please

ganeshie8 (ganeshie8):

when the 2 resistors are connected in parallel ideally we expect to see 100 * 200/300 = 66.67 ohm. but since these are not ideal, first resistor can change +- 3, that is, it can be anywhere between [97, 103] second resistor can change +- 4, that is, it can be anywhere between [196, 204] so overall, the parallel connection can see a resistance of : [97||196 103||204] = [64.887, 68.443] = -1.78, +1.78 = +-1.78 so thats an absolute error of +-1.78 and a % error of 1.78/66.67*100 = 2.67 %

OpenStudy (anonymous):

Thank you, thank you, THANK A LOT!! ^_^ :D

ganeshie8 (ganeshie8):

np hope u understood the working :)

OpenStudy (anonymous):

100 * 200/300 = 66.67 ohm. how it just come?

OpenStudy (anonymous):

okayy.. GOT IT.. :D

ganeshie8 (ganeshie8):

good to hear :)